you are looking at a map that uses cartesian coordinates. Your school is exactly halfway between your house in your friends house. You live at (-2, 3) and the school is at (7, 10). What are the coordinates of where your friend lives?


1
Expert's answer
2021-08-27T12:51:05-0400

Midpoint formula:

"x=\frac{x_1+x_2}{2};\space y=\frac{y_1+y_2}{2}."

"(x,y)" are the coordinates of the school;

"(x_1,y_1)" are the coordinates of my home;

"(x_2,y_2)" are the coordinates of my friend's house;

"x_2=2x-x_1=14+2=16;"

"y_2=2y-y_1=20-3=17;"

Answer:

My freand lives at "(16,17)."